OpenStudy (anonymous):

Octavian became the emperor of Rome in 27 B.C. He ruled until his death in A.D. 14. How long did he rule as emperor? (Hint: Remember there was no A.D. 0.) 40 years 41 years 13 years 12 years

OpenStudy (anonymous):

41 yrs

jimthompson5910 (jim_thompson5910):

The years that end with BC count down as time progresses forward. So if someone was born in 27 BC, then they'd be 1 year old in 27-1 = 26 BC So 27 years pass going from 27 BC to 0 BC (since 27-27 = 0) Now add on 14 since you're going from 1 AD to 14 AD to get 27 + 14 = 41 So Octavian ruled for 41 years.